# Field transformations in functional integral, effective action and functional flow equations

### v2

@@ Conclusions
The present paper addresses a systematic discussion of field transformations, starting from variables in the functional integral and proceeding to transformations of field variables in the effective action and for flow equations for associated scale-dependent effective average actions. In order to illustrate the conceptual discussion with a practical case we have focused often on a simple scalar model with a field-dependent kinetic term.
For the field transformations in functional flow equations we distinguished between two different concepts – the macroscopic and the microscopic field transformations. The macroscopic field transformations are transformations of the variables in terms of which a given effective average action is expressed. They do not change the effective average action and the associated field equations and macroscopic correlation functions. These quantities are only expressed in terms of different “field coordinates”. As familiar from differential equations, a suitable choice of variables can be very helpful to find (approximate) solutions. In quantum field theory the transformation to renormalized fields can absorb part of the microscopic information which does not affect the macroscopic observations. The use of dimensionless renormalized fields reveals the possible existence of scaling solutions. Selecting suitable composite fields can be of great help to make the relevant macroscopic degrees of freedom easily visible.
@@
For linear microscopic field transformations the different source terms simply result in a linear macroscopic field transformation. For linear transformations the macroscopic and microscopic field transformations are equivalent. This does no longer hold for non-linear transformations. Macroscopic correlation functions are defined by functional derivatives of the effective action. In the non-linear case there is no simple map between the macroscopic correlation functions with or without the microscopic field transformation. Microscopic field transformations typically loose the direct contact to the microscopic correlation functions of the original model. This contact is maintained for the macroscopic field transformations as long as the transformation to canonical fields is known explicitly.

### v3

@@ Discussion
For the field transformations in functional flow equations we distinguish between two different concepts – the macroscopic and the microscopic field transformations. The macroscopic field transformations are transformations of the variables in terms of which a given effective average action is expressed. They do not change the effective average action and the associated field equations and macroscopic correlation functions. These quantities are only expressed in terms of different “field coordinates”. As familiar from differential equations, a suitable choice of variables can be very helpful to find (approximate) solutions. In quantum field theory the transformation to renormalized fields can absorb part of the microscopic information which does not affect the macroscopic observations. The use of dimensionless renormalized fields reveals the possible existence of scaling solutions. Selecting suitable fields can be of great help to make the relevant macroscopic degrees of freedom easily visible.
@@
For linear microscopic field transformations the different source terms simply result in a linear macroscopic field transformation. For linear transformations the macroscopic and microscopic field transformations are equivalent. This does no longer hold for non-linear transformations. Macroscopic correlation functions are defined by functional derivatives of the effective action. In the non-linear case there is no longer a direct map between the macroscopic correlation functions and microscopic correlations, neither for the original canonical field nor for composite microscopic fields. This contact is can be reconstructed for the macroscopic field transformations as long as the transformation to canonical fields is known explicitly. Microscopic field transformations of the average effective action typically loose the direct contact to the microscopic correlation functions for the canonical fields of the original model. This is also no direct relation to correlations of microscopic composite fields if the relation between the canonical and composite fields is not known for a given RG-kernel.